Circular No. 7051

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES AND RINGS OF NEPTUNE C. Roddier, F. Roddier, J. E. Graves, O. Guyon, and M. J. North-

cott, Institute for Astronomy, University of Hawaii (UH); and B. Sicardy, Paris-Meudon Observatory, report: “Infrared (1.72-μm) images of Neptune taken on July 6 with the UH adaptive optics system (Hokupa’a) mounted on the Canada-France-Hawaii 3.6-m telescope have now been thoroughly processed. In addition to the observations of Neptune VI (Galatea), VII (Larissa), and VIII (Proteus) reported on IAUC 6987, Neptune V (Despina) and Neptune’s ring arcs have also been detected. Despina, which was far from maximum elongation, moved substantially during the 600-s exposure time, but its average position is consistent within ± 1^{o} of that estimated from Voyager 2 data in 1989 (as for Proteus and Larissa). Each of the three brightest arcs Liberté, Egalité, and Fraternité are identified. The fainter leading arc Courage cannot be confirmed. The positions of the main arcs are close to that given by the second orbital solution of Nicholson et al. (1995, Icarus 113, 295), which corresponds to a mean motion of 820^{o}.1118 day^{−1}. Photometric profiles of the arcs have been obtained and compared with those of Porco (1991, Science 253, 995), allowing for the blur due to our 600-s integration time. For the two trailing arcs (Fraternité and Egalité), the fit is very good. It shows that the arcs are 4^{o} ± 2^{o} ahead of the position estimated from Nicholson’s second solution. For Liberté the uncertainty is larger due to noise. Four 200-s exposures made in the J band were also analyzed. The arcs were not detected. However, the data confirm that Galatea is 5^{o} ± 1^{o} ahead of the nominal predicted position (IAUC 6987). In these data, Despina is closer to maximum elongation. Its image is sharper and its position is again consistent within ± 1^{o} with that predicted from Voyager data. We note that the uncertainties in the orbital parameters derived from Voyager data by Owen et al. (1991, A.J. 101, 1511) lead to the following standard errors on the current satellite longitudes: ± 2^{o}.9 for Proteus, ± 5^{o}.2 for Larissa, ± 9^{o}.1 for Despina, and ± 8^{o}.1 for Galatea. The estimated errors are believed to be conservative by typically a factor of two (ibid.). Further analysis is thus required to see if the Galatea O–C is significant or a mere effect of the accumulation of error between 1989 and 1998.”

COMET P/1998 U3 (JÄGER) Visual m_{1} estimates by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): Oct. 30.84 UT, 12.0; Nov. 11.80, 11.2; 12.82, 11.3.

1998 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green